The length of a rectangle is five less than two times the width. If the perimeter is 50 inches, find the length and the width.
Answer by mananth(16946)

width be x
length 2x-5
..
Perimeter P = 2(L+W)
2*(x+2x-5)
=2(3x-5)
=6x-10 = 50
6x= 50+10
6x=60
x= 60/6
x= 10
length = 2x-5
= 2*10-5
=15